OpenStudy (juliebeans):

If f(x) varies directly with x and f(x) = 3 when x = 12, then what is the value of f(x) when x = 40? 480 160 10 4

OpenStudy (thatonegirl_):

The formula for direct variation is y(or f(x)=kx. So plug in what you know into that equation and solve for k. Now that you know k, plug in what you know into f(x)=kx and you can solve for f(x).

OpenStudy (juliebeans):

So f(x)=12(40)

OpenStudy (thatonegirl_):

3=k(12) 3/12=k 1/4=k f(x)=1/4(40) f(x)=10

OpenStudy (thatonegirl_):

First you solve for k. Then with the second set of numbers find f(x) now that you know what k is

OpenStudy (thatonegirl_):

Do you get it?

OpenStudy (juliebeans):

Kinda of thanks

OpenStudy (juliebeans):

so the answer is d?"

OpenStudy (thatonegirl_):

It would be 10

OpenStudy (thatonegirl_):

Do u have another practice problem?

OpenStudy (juliebeans):

Yes :P

OpenStudy (juliebeans):

Given the function T(z) = 7 – z, find T(–4). –11 11 -3 3

OpenStudy (thatonegirl_):

Is that \[t(z)=7-z\]

OpenStudy (thatonegirl_):

Since it gives you T(z) is -4, you would just plug that in. -4=7-z Subtract the 7 -11=z

OpenStudy (juliebeans):

oh

OpenStudy (thatonegirl_):

Is that what you were thinking?

OpenStudy (juliebeans):

Yes